# KPNDepth: Depth Estimation of Lane Images under Complex Rainy Environment

### v1

@@ Abstract
With the development of deep neural network generative models in recent years, significant progress has been made in the research of depth estimation in lane scenes. However, current research achievements are mainly focused on clear daytime scenarios. In complex rainy environments, the influence of rain streaks and local fog effects often leads to erroneous increases in the overall depth estimation values in images. Moreover, these natural factors can introduce disturbances to the accurate prediction of depth boundaries in images. In this paper, we investigate lane depth estimation in complex rainy environments. Based on the concept of convolutional kernel prediction, we propose a dual-layer pixel-wise convolutional kernel prediction network trained on offline data. By predicting two sets of independent convolutional kernels for the target image, we restore the depth information loss caused by complex environmental factors and address the issue of rain streak artifacts generated by a single convolutional kernel set. Furthermore, considering the lack of real rainy lane data currently available, we introduce an image synthesis algorithm, RCFLane, which comprehensively considers the darkening of the environment due to rainfall and local fog effects. We create a synthetic dataset containing 820 experimental images, which we refer to as RainKITTI, on the commonly used depth estimation dataset KITTI. Extensive experiments demonstrate that our proposed depth estimation framework achieves favorable results in highly complex lane rainy environments.
@@ I Introduction
Depth estimation is one of the key tasks in the field of computer vision, aiming to infer distance information from each pixel in an image to the camera. Based on this depth information, the geometric layout of a three-dimensional scene can be quickly understood, providing crucial visual information support for areas such as autonomous driving and intelligent transportation systems. Traditional depth estimation often relies on SLAM methods [1], which, in the absence of prior knowledge, use sensors (such as high-precision depth cameras, LiDAR, etc.) for self-localization and environmental modeling. However, high-precision depth cameras and LiDAR systems often come with high equipment costs. To address this issue, in recent years, with the advancement of deep neural networks, estimating pixel-level dense depth from a single image captured by a regular camera is becoming feasible [2]. As research progresses, various types of deep learning-based monocular depth estimation frameworks have been proposed, including supervised learning methods that utilize depth information from 3D sensor scans as training ground truth [3], as well as self-supervised and unsupervised methods based on binocular epipolar geometric constraints [4] and monocular frame sequence pose estimation [5], separately. These methods have garnered widespread attention in recent years due to their low cost and ability to generalize across various scenarios.
However, current research on monocular depth estimation has long faced the challenge of complex weather conditions. Additional natural factors such as occlusion by rain streaks and local fog effects [6] can lead to errors and loss of depth information in images. Moreover, the KITTI dataset [7], which is widely used for monocular depth estimation, only contains road images during clear daytime and lacks driving data on rainy days and nights. Therefore, existing studies [8][9] rely on synthetic data for training. However, existing research indicates that unsupervised approaches using geometric constraints instead of ground truth face an inevitable dilemma when confronted with augmented datasets. Existing depth estimation algorithms, especially self-supervised and unsupervised methods, are based on the assumption of photometric consistency [10], which is generally satisfied in daytime images where the sun serves as the primary light source. However, in complex rainy conditions, multiple light sources such as streetlights, car lights, and high-reflectance surfaces due to rainwater violate inter-frame photometric consistency, leading to significant errors in photometric loss calculation and severely impacting the estimation of depth information in images.
@@
When introducing complex rainy lane images, we diverge from current mainstream approaches [11][12] that directly input synthetic images alongside original images into the depth estimation framework for self-supervision. Instead, we decompose depth estimation in complex environments into two subtasks: image reconstruction and depth estimation. By constructing an efficient and accurate image reconstruction network, DLKPN, trained on offline data, we aim to restore depth information in synthetic images fed into the depth estimation network. This approach aims to reduce the reliance of the proposed Added Depth Estimation Model of DLKPN (KPNDepth) framework on the assumptions of photometric consistency and static scene conditions in input images. Our main contributions are as follows:
  • Taking into account the environmental darkening and local fog effects caused by rain, we have proposed an algorithm called RCFLane for synthesizing complex rainy lane images. Based on this algorithm, we have synthesized RainKITTI using the KITTI dataset [7], and we plan to release this dataset.
  • Building upon the concept of convolutional kernel prediction [13], we have developed the image reconstruction network DLKPN to address rain streak artifacts in complex rainy lane image reconstruction. We have achieved efficient registration of input images on RainKITTI.
  • ·We have tested our model on RainKITTI, achieving more effective image reconstruction compared to existing methods [14][15]. Furthermore, the proposed KPNDepth framework has shown more accurate depth estimation results compared to baseline models [16].
@@ II PROPOSED METHOD
In this section, we will introduce the complete KPNDepth framework, the RCFLane algorithm for synthesizing real rainy lane images, and the DLKPN as the core of this framework.
@@ II-A RCFLane: Building Real Lane Data in Rainy Cloudy and Foggy Days
Due to the effect of rain on ambient lighting and camera exposure, it is almost impossible to take a realistic set of photos from the same location, leaving the only difference between the two photos being the presence or absence of rain streaks. Most existing commonly used datasets add 2D rain streaks to original image data. However, for the task we are studying, these datasets only consider the factor of rain streaks when synthesizing rainy day datasets, while ignoring the darkening of the environment and local fog effects caused by rainfall. These factors are crucial in our depth estimation task and cannot be overlooked. Although the RainCityscapes dataset proposed in [17] introduces fog when synthesizing data, it focuses on urban landscapes rather than lane data, which does not align with the training purpose of our model. Therefore, this paper will start with the commonly used KITTI dataset for lane depth estimation, aiming to simulate realistic rain characteristics as much as possible on it to construct subsequent training and testing data.
@@ II-A1 Rain Layer
In the rain operation of RCFLane, we rely on a Gaussian blur kernel. By introducing three parameters - rotation center coordinates, rotation angle, and scale factor - we construct an affine transformation matrix to perform image rain filtering by rotating and generating a Gaussian blur kernel based on the unit focus matrix.
It is important to note that the rain layer constructed using Gaussian blur has a black background. We only want to overlay the white rain lines, so in Eq.(1) we set the parameter $\alpha$
@@ II-A2 Mask Layer
In reality, the impact of rain on images is not only manifested in the obscuration caused by rain streaks but also in the deterioration of ambient lighting conditions. To reflect the changes in lighting conditions due to rainy days, this paper introduces a mask layer on top of the previous $O_{1}(x)$ in Eq.(2). The mask layer can be understood as a fully black curtain, weighted by a parameter $\gamma$ set by humans, to simulate environmental lighting from day to night. This allows for a more realistic simulation of lighting conditions on rainy days, making the synthetic data closer to real rainy day lighting situations.
@@ II-A3 Transmission Layer
The impact of rain curtains on images presents a certain fog-like characteristic on the images visually due to the scene depth. This fog-like characteristic becomes more pronounced as the distance from the camera increases.
In this paper, the addition of fog in Eq.(3) is based on the standard optical model. In imaging under hazy weather conditions, the reduction of reflected energy causes light transmission attenuation, and atmospheric light scatters around, enhancing the surrounding brightness in foggy conditions. Therefore, the pixel value at a certain point in the foggy image consists of two parts
@@
The parameter $S$ represents the fog scale, where there is a negative correlation between $d(x)$ (the distance from a pixel point to the center point) and the fog scale. The reason why the RCFLane algorithm fixes the fogging center at the image center instead of incorporating image depth information into fog synthesis, as suggested in [6], is twofold. Firstly, the method proposed in [6] demands precise image depth values, we hope to explore a less restrictive lane rainy day image synthesis method. Secondly, and more importantly, for lane data, the center of the image often represents the end of the road extension, that is, the farthest distance from the camera. Based on the nature of rain curtain fog being distance-dependent, we can infer that the fog thickness reaches its maximum at the image center and decreases towards the periphery. Therefore, fixing the fogging center at the image center is deemed reasonable.
@@ II-B Dual-layer Pixel-wise Convolution Kernel Prediction Network
Whether it is rain or the decrease in illumination caused by rain curtains and local fog effects can be considered as image degradation. We hope to effectively restore these degradations before inputting the images into the depth estimation network, thereby addressing the depth increase and local pseudo-artifacts caused by rain curtains. Additionally, due to downstream requirements, today’s depth estimation tasks often have high requirements for real-time prediction, which reminds us to pay attention to the inference time required for the reconstruction algorithm.
Inspired by recent research on image restoration and rain removal [13][18][19], we aim to establish a kernel prediction network for rain removal, predicting convolutional kernels instead of directly predicting the pixel values of the target pixels. It has been proposed in [13][18][19] that predicting pixel-wise convolutional kernels for image restoration offers advantages in terms of convergence speed and inference speed, aligning well with the objectives of our study.
@@
However, through experiments, we found that the degradation in $\hat{O}(x)$ compared to $O(x)$ is not solely due to rain streaks. Introducing a single KPN network with pixel-wise convolutional kernels predicted for each pixel cannot fully explain this degradation. As a result, the reconstructed image still retains some degradation information, manifested notably as rain streak artifacts. Fig. 3 specifically demonstrates this issue using predictions on RainKITTI data. Detailed experimental results will be presented in Section IV. Therefore, we have constructed a Dual-layer Pixel-wise Convolution Kernel Prediction Network (DLKPN). This network is trained offline on two independent sets of data. One layer is dedicated to recovering most of the degradation information, while the other layer focuses primarily on removing the artifacts caused by rain streaks. For each pixel $p$ in $\hat{O}(x)$, the network will predict two sets of independent pixel-wise convolution kernels $K_{p1}$ and $K_{p2}$
@@ II-C Added Depth Estimation Model of DLKPN
The DLKPN model proposed in this paper addresses the reconstruction of rainy lane images, with the advantage of fast inference speed, aligning well with the real-time requirements of depth estimation tasks. We will offline train the required pre-trained DLKPN model on the RainKITTI dataset and integrate it with traditional depth estimation models to enhance the accuracy of depth estimation in rainy environments. For ease of description, we will refer to this framework as KPNDepth.
As a baseline, we have chosen LapDepth[16], a model that utilizes Laplacian pyramids as part of the decoder structure. It takes the encoded features and feeds them into different streams, with each stream responsible for decoding depth residuals. By decomposing the Laplacian pyramid to define depth residuals and gradually combining the corresponding outputs from coarse to fine, it reconstructs the final depth image. The overall structure of the KPNDepth framework can be referenced in Fig. 5.
@@ III-A Datasets
Since this paper focuses on rainy lane depth estimation and there is currently a lack of publicly available rainy datasets specialized in lane information, we started from the commonly used public depth estimation dataset KITTI[7] and used the RCFLane algorithm to synthesize the RainKITTI dataset used in this study. The KITTI dataset is one of the most commonly used benchmark datasets for computer vision algorithms in autonomous driving scenarios internationally, containing real image data from urban, rural, and highway scenes. We randomly selected 820 images from the left camera images of the $2011\_09\_28$ and $2011\_09\_29$ sequences in KITTI and synthesized RainKITTI through rain image synthesis. The resolution of the synthesized images will be kept consistent with the original images (1238 x 374). For each synthesized image, we retained its ground truth depth, which will be used to evaluate the performance of our depth estimation model.
  • Training Sets:For the RainKITTI dataset, we have partitioned 715 images as the training set to train the DLKPN network structure.
  • Testing Sets:In this experiment, we have set aside 105 images from RainKITTI as the testing set. This testing set will be used to evaluate the performance of the dual-layer KPN network reconstruction. For the reconstructed images, we will assess the specific performance of the KPNDepth framework in rainy lane depth estimation based on their real depth information and depth network predictions.
@@ III-B Experimental Setup
When using the RCFLane algorithm to synthesize rainy images, the weights introduced for the rain layer and mask layer are denoted as follows: weight $\beta$ for the rain layer, weight $\gamma$ for the mask layer, attenuation coefficient $\lambda$ and atmospheric light intensity $A$ are set to $\beta=1.0$, $\gamma=0.2$, $\lambda=0.025$ and $A=0.5$, respectively. Faced with various atmospheric lighting conditions and different levels of rainfall leading to diverse weather scenarios, modifying the corresponding parameters allows for the synthesis of rainy lane images in the target environment. This capability holds practical significance for depth estimation in complex rainy environments.
During the training of DLKPN, for the first layer KPN network, we set the batch size to 16 and trained for a total of 1000 epochs. We used the inference results of the pre-trained network from the first layer as the training input for the second layer KPN, focusing on removing the rain streak artifacts introduced by the first layer model. With the same batch size, we recorded the network’s performance on the test set in terms of PSNR and SSIM every 100 epochs. After experimentation, we selected the model trained for 400 epochs as the pre-trained network for the second layer of KPN.
III-C Results
@@
III-C1 Image Reconstruction on RainKITTI
The relevant experimental results are shown in Table I. Since the DLKPN method proposed in this paper is based on our construction of the RainKITTI dataset for training and testing, it is not possible to obtain usable prior data from the literature. Moreover, related studies on the fog effects of rain curtains in rain image reconstruction, represented by [17], have not been open-sourced. To evaluate the actual effectiveness of the DLKPN network, we selected the state-of-the-art (SOTA) model DRSformer[14] and the classic lightweight rain removal model PreNet[15] in the field of rain image restoration. After training the above models on the RainKITTI dataset, we compared them with DLKPN in terms of Quantitative Evaluation Metric and inference latency.
The models PreNet, PReNet_r, PRN, and PRN_r in the table are all model architectures proposed in [15], differing in whether recursive layers and intra-level recursive calculations are introduced. The inference latency in Table I is represented by Times(ms), reflecting the inference time of a single image.
From the results, it can be seen that the DLKPN network achieved better SSIM and PSNR on the RainKITTI dataset than the current SOTA method [14] because it considers the darkening of the environment and local fog effects caused by rain when removing rain streaks. By comparing the restored images output by each model in Table 1, it can be observed that the reason for the lower SSIM and PSNR of methods like DRSformer is that these methods still retain some mask and fog information in the reconstructed images. In terms of inference latency, the single-image inference time of DLKPN is 152.89ms. Although there is still a gap compared to the lightweight model PRN_r, it has improved the speed by 4.94 times compared to the SOTA model DRSformer. Considering the SSIM and PSNR performance of the overall model, we believe that the inference latency of DLKPN is acceptable.
III-C2 Depth Estimation of Lane in Rainy Days
@@
III-C3 Ablation Study
In the establishment of the DLKPN network, we followed the idea of predicting convolutional kernels instead of directly predicting the pixel values of reconstructed images, as seen in [13][18][19]. Building upon the single KPN network model in [18], we introduced another set of independent pixel-wise convolutional kernels focused on removing rain streak artifacts.
To analyze the impact of the newly introduced kernel set on image reconstruction, we designed a series of ablations to evaluate our method, with quantitative analysis results presented in Table III. The experiments demonstrate that the two sets of independent pixel-wise convolutional kernels introduced by DLKPN significantly enhance the quality of reconstructed images, providing better robustness on lane images in complex weather conditions. However, it is important to note that due to the additional convolutional computation introduced by this network, the inference time for a single image will correspondingly increase.

### v2

@@ Abstract
Recent advancements in deep neural networks have improved depth estimation in clear, daytime driving scenarios. However, existing methods struggle with rainy conditions due to rain streaks and fog, which distort depth estimation. This paper introduces a novel dual-layer convolutional kernel prediction network for lane depth estimation in rainy environments. It predicts two sets of kernels to mitigate depth loss and rain streak artifacts. To address the scarcity of real rainy lane data, an image synthesis algorithm, RCFLane, is presented, creating a synthetic dataset called RainKITTI. Experiments show the framework’s effectiveness in complex rainy conditions.
@@ I Introduction
Depth estimation is crucial in computer vision for determining the distance of each pixel to the camera, which is key to understanding 3D scene geometry and essential for applications like autonomous driving and intelligent transportation systems. Traditional approaches, such as SLAM [1], have relied on expensive high-precision sensors like depth cameras and LiDAR for self-localization and environmental mapping. However, advances in deep neural networks have made it possible to perform dense depth estimation from a single image using a regular camera [2]. Various frameworks for monocular depth estimation have emerged, including supervised learning with 3D sensor data [3], and self-supervised and unsupervised methods leveraging binocular geometry [4] and monocular sequence pose estimation [5]. These methods are gaining popularity due to their cost-effectiveness and versatility across different scenarios
However, monocular depth estimation faces significant challenges in complex weather conditions, such as rain streaks and local fog effects [6], which can lead to errors and loss of depth information in images. The widely used KITTI dataset [7] for monocular depth estimation contains only clear daytime road images and lacks data for rainy days and nights. Consequently, existing studies [8][9] depend on synthetic data for training. However, due to the large differences in weather, lighting, etc. between daytime lane images and rainy or night lane images, when the existing monocular depth estimation model is applied to complex environment images, there may be significant domain deviations, resulting in a notable reduction in depth estimation performance, as indicated in [14]. Therefore, existing monocular estimation research focuses on how to build a domain adaptation framework to transfer models trained on daytime images to nighttime or rainy day data sets. At the same time, unsupervised approaches that use geometric constraints instead of ground truth encounter difficulties with augmented datasets, as they are predicated on the assumption of photometric consistency [10], which is often violated in rainy conditions due to the presence of multiple light sources like streetlights, car lights, and reflective rainwater surfaces. This violation results in significant errors in photometric loss calculation and severely impacts the estimation of depth information in images. Moreover, existing rainy day depth estimation solutions still regard depth estimation in complex environments as an independent task, which makes the depth model increasingly complex.
@@
When introducing complex rainy lane images, we depart from mainstream methods [11][12] that integrate synthetic and original images for self-supervised depth estimation. We segment the task into image reconstruction and depth estimation. Our efficient and precise image reconstruction network, DLKPN, trained offline, is designed to refine synthetic images before they are processed by the depth estimation network, thereby lessening the KPNDepth framework’s reliance on photometric consistency and static scene assumptions.
  • We introduce RCFLane, an algorithm accounting for environmental darkening and local fog effects in rain, to synthesize complex rainy lane images. Utilizing this, we created the RainKITTI dataset [7], which we intend to release.
  • Leveraging convolutional kernel prediction [13], we developed DLKPN to counteract rain streak artifacts in image reconstruction on RainKITTI, achieving effective image registration.
  • Our model, tested on RainKITTI, outperforms existing methods [14][15] in image reconstruction. Moreover, the KPNDepth framework demonstrates superior accuracy in depth estimation compared to baseline models [16].
@@ II-A RCFLane: Building Real Lane Data in Rainy Cloudy and Foggy Days
Rain’s impact on lighting and camera exposure precludes capturing realistic photos with mere variations in rain streak presence. While existing datasets typically add 2D streaks, they overlook environmental darkening and fog, essential for depth estimation tasks. The RainCityscapes dataset[17], though fog-inclusive, targets urban landscapes, not lane data, diverging from our model’s training objectives. Consequently, this paper adopts the KITTI dataset to simulate realistic rain effects for constructing lane depth estimation data.
@@ II-A1 Rain Layer
In RCFLane’s rain simulation, an affine transformation matrix is constructed using Gaussian blur with parameters for rotation center, angle, and scale, enabling image rain effects. Notably, the Gaussian-generated rain layer has a black backdrop, necessitating the overlay of white rain streaks, which is achieved by adjusting parameter $\alpha$ in Eq.(1)
@@ II-A2 Mask Layer
In reality, the impact of rain on images is not only manifested in the obscuration caused by rain streaks but also in the deterioration of ambient lighting conditions. To reflect the changes in lighting conditions due to rainy days, this paper introduces a mask layer on top of the previous $O_{1}(x)$ in Eq.(2). The mask layer is weighted by a human-set parameter $\gamma$, to mimic the transition from day to night lighting, enhancing the realism of synthetic data for rainy conditions.
@@ II-A3 Transmission Layer
Rain-induced fog visually imparts a characteristic to images that intensifies with increasing distance from the camera. The paper’s fog addition in Eq.(3) adheres to the optical model, where light transmission is attenuated by reduced reflected energy, and atmospheric light scattering boosts brightness in fog. Therefore, the pixel value at a certain point in the foggy image consists of two parts
@@
The parameter $S$, inversely proportional to the distance $d(x)$ from the image center, represents fog scale. The RCFLane algorithm employs a central fogging point for rainy lane image synthesis, diverging from depth-guided methods like those in [6]. This choice is motivated by two factors: it circumvents the need for accurate depth data, enhancing flexibility, and it aligns with the center’s significance in lane imagery, representing the farthest point from the camera. Given the distance-dependent nature of rain-induced fog, the algorithm rationally assumes peak fog density at the center, tapering off at the periphery, thus justifying the focus on central fog.
@@ II-B Dual-layer Pixel-wise Convolution Kernel Prediction Network
Rain and the resulting illumination reduction and local fog are forms of image degradation we aim to restore before depth estimation to mitigate rain-induced depth increases and artifacts. The need for real-time depth prediction in modern tasks also demands efficient inference times for our restoration algorithm.
Drawing from recent advances in image restoration and rain removal [13][18][19], we propose a kernel prediction network that predicts convolutional kernels rather than direct pixel values for the target pixels. The approach, supported by [13][18][19], is advantageous for faster convergence and inference, meeting our study’s goals.
@@
Experiments revealed that degradation in $\hat{O}(x)$ relative to $O(x)$ is not just from rain streaks, and a single KPN network with pixel-wise kernels is insufficient for complete degradation explanation, leaving residual artifacts. Fig. 3 illustrates this on RainKITTI data, with full results in Section IV. To address this, we developed the Dual-layer Pixel-wise Convolution Kernel Prediction Network (DLKPN), trained on distinct datasets to recover degradation and remove rain streak artifacts. For pixel $p$ in $\hat{O}(x)$, DLKPN predicts two independent kernel sets $K_{p1}$ and $K_{p2}$:
@@ II-C Added Depth Estimation Model of DLKPN
The DLKPN model, optimized for rapid inference, is proposed for rainy lane image reconstruction to meet real-time depth estimation demands. We will train the DLKPN model offline using the RainKITTI dataset and combine it with conventional models to improve rainy-day depth estimation accuracy, termed as KPNDepth. We chose LapDepth[16] as the baseline, utilizing Laplacian pyramids to decode depth residuals from multi-stream features, refining the depth image progressively. Refer to Fig. 5 for the KPNDepth framework’s architecture.
@@ III-A Datasets
This paper tackles the challenge of rainy lane depth estimation, an area where specialized datasets are limited. To address this, we leveraged the KITTI dataset [7], a well-established benchmark for autonomous driving vision algorithms that includes diverse urban, rural, and highway scenes. Utilizing the RCFLane algorithm, we synthesized the RainKITTI dataset from the $82011_{0}9_{2}$ 8 and $92011_{0}9_{2}$ 9 sequences, carefully selecting 820 images while preserving their original resolution of 1238 x 374 pixels. This dataset was then split into training and test sets, with the ground truth depth information retained for thorough model evaluation.
@@ III-B Experimental Setup
III-C Metrics
Rainy Day Image Reconstruction Metrics: We use peak signal-to-noise ratio (PSNR) and structural similarity (SSIM), which are commonly used in the field of image reconstruction, as quantitative evaluation metrics for all datasets. Here, the larger the PSNR and SSIM are, the better the rain removal effect is. Times (ms) represents the processing time of a single image.
Rainy Day Lane Depth Estimation Metrics: For the quantitative evaluation of the depth estimation effect, we use six indicators introduced by Eigen et al. [20] Abs Rel, Sq Rel, RMSE, RMSE log, Accuracy and log10, which have been widely used and verified in the field of depth estimation.
III-D Results
@@
III-D1 Image Reconstruction on RainKITTI
The experimental results, presented in Table I, demonstrate the performance of the DLKPN method on the RainKITTI dataset, which lacks comparable prior data due to its novel construction. The absence of open-sourced studies, such as [17], on rain curtain fog effects further limits direct comparisons. To assess DLKPN’s efficacy, we benchmarked it against the SOTA model DRSformer[14] and the lightweight PreNet[15] on both Quantitative Evaluation Metrics and inference latency.
The table includes variations of PreNet[15], which differ by the presence of recursive layers. Inference latency, measured in Times(ms), indicates the processing time for a single image. DLKPN outperformed [14] in SSIM and PSNR by accounting for rain-induced environmental darkening and fog, evident in the visual quality of restored images. Despite a higher latency than the lightweight PRN_r, DLKPN’s 152.89ms single-image inference time is 4.94 times faster than DRSformer, making it a balanced choice given its superior SSIM and PSNR.
III-D2 Depth Estimation of Lane in Rainy Days
@@
III-D3 Ablation Study
In developing the DLKPN network, we adopted the strategy of kernel prediction over direct pixel value prediction, as per [13][18][19]. Expanding on [18]’s single KPN model, we added an independent set of kernels to eliminate rain streak artifacts.
To assess the contribution of the new kernels, we conducted ablation studies, with results in Table III. These show that DLKPN’s dual kernels notably improve reconstruction quality and robustness on lane images in adverse weather.
The DLKPN model, while outperforming the baseline in computational accuracy due to its enhanced convolutional operations, incurs longer computation times for single image processing. In scenarios with simple lighting, this extra computational load is deemed unnecessary. Given the DLKPN framework’s dual-module structure, we propose a preliminary assessment of the lane environment before inference. For environments with minimal rain streaks and fog, computation is restricted to the single-layer KPN network, reducing unnecessary latency and improving model efficiency in real-world applications.